# Hestia: Voxel-Face-Aware Hierarchical Next-Best-View Acquisition for Efficient 3D Reconstruction

### v2

@@ 1 Introduction
Multiview-based 3D scene reconstruction [schoenberger2016sfm, yao2018mvsnet, xie2019pix2vox, murez2020atlas, wang2021multi, sayed2022simplerecon, xu2023unifying, wang2024dust3r, duisterhof2024mast3r, pan2024glomap] and novel view synthesis [mildenhall2020nerf, yu2021pixelnerf, yu2022plenoxels, mueller2022instant, zheng2024gpsgaussian, gslrm2024, charatan2024pixelsplat, sabour2024spotlesssplats, flynn2024quark, chen2025mvsplat] have been central topics in computer vision. These methods leverage multiview information from images to reconstruct high-fidelity scenes or synthesize photorealistic novel views. However, data acquisition remains a bottleneck. Most data is collected manually, which is time-consuming and labor-intensive, or follows preplanned camera trajectories that do not generalize well to diverse objects.
@@ 2 Related Work
Scene-specific next-best-view planners. Next-best-view planners have demonstrated promising results in active 3D reconstruction by predicting the optimal viewpoint for data capture based on the current state. Traditional approaches [monica2018contour, zha1997next, liu2018object, hardouin2020surface, hardouin2020next, dai2020fast, zhou2021fuel] rely on hand-crafted rules to determine the next-best viewpoint. For instance, the method [zha1997next] selected the next-best viewpoint by maximizing a rating function favoring smooth regions, which may overlook fine-grained object details. Instead, the methods [monica2018contour, hardouin2020surface] collected data along boundaries between seen and unseen surfaces to capture finer details, but still require handcrafted parameter tuning for each scene. Another approach [liu2018object] scanned segmented objects sequentially using a predefined object database, reducing the need for handcrafted tuning, but its performance degrades in cluttered environments due to inaccurate object matching. Recent advances in deep learning and increased computational power have given rise to learning-based methods [lin2022active, pan2022activenerf, lee2022uncertainty, zhan2022activermap, jin2023neu, sunderhauf2023density, chen2024gennbv, peralta2020next, ran2023neurar, zeng2022efficient, tao2022seer]. Some studies [lin2022active, sunderhauf2023density] used NeRF [mildenhall2020nerf] ensembles to estimate uncertainty via model disagreement for viewpoint selection, resulting in linearly increasing computational overhead. Other works [pan2022activenerf, ran2023neurar, zeng2022efficient] avoid the computational overhead by incorporating Bayesian-based NeRF [shen2021stochastic, martin2021nerf] to estimate uncertainty for viewpoint selection. Meanwhile, other approaches [lee2022uncertainty, zhan2022activermap] defined the next-best viewpoint as the viewpoint that maximizes the entropy of the density field along the camera rays. Although these methods have shown outstanding performance in collecting data, all require sampling candidate viewpoints. In addition, their reliance on online learning makes them less suitable for real-time applications.
@@
Generalizable next-best-view planners. Unlike the aforementioned online-learning approaches, the generalizable methods [jin2023neu, chen2024gennbv, peralta2020next] avoid the training process for new scenes, thereby enabling faster next-best-view selection. Prediction time is important for real-world tasks where a robot (e.g., a drone) may run out of battery within a few minutes. Among generalizable methods, prior work [jin2023neu] proposed a Bayesian-based NeRF that selects the next-best viewpoint to maximize view variance without additional training. However, it still requires candidate viewpoint sampling, resulting in performance unsuitable for real-time applications. Instead, another line of generalizable next-best-view approaches [chen2024gennbv, peralta2020next] utilized reinforcement learning to learn a next-best-view planner to bypass the need for sampling candidate viewpoints. Prior work [peralta2020next] proposed learning a 3-DoF next-best-view policy using a series of grayscale images as observations. Subsequently, prior work [chen2024gennbv] improved upon this method by incorporating occupancy grids into the observations, which provide explicit geometric information. This enhancement enabled the development of a 5-DoF next-best-view planner, achieving an outstanding coverage ratio for unknown scenes.
@@
As shown in Tab. S4, compared to the scene-specific methods, Hestia does not require candidate viewpoint sampling or inference-time optimization, thereby enabling more flexible viewpoint prediction. Compared to the generalizable method [chen2024gennbv], Hestia treats voxels as cubes rather than points, enabling more comprehensive capture. In addition, Hestia adopts a close-greedy scheme to mitigate spurious correlations, introduces a hierarchical structure to model the action space, and uses a more diverse dataset to maintain robust performance across varying object shapes and positions. Notably, Hestia’s hierarchical structure addresses the challenge of high-dimensional continuous action search in RL-based generalizable next-best-view planning. This differs from traditional methods [monica2018contour, hardouin2020surface, dai2020fast, zhou2021fuel], which use hierarchical structures to traverse frontiers or optimize tour efficiency.
@@ 3.1 RL Problem Definition
where $\mathbbm{1}(\cdot)$ is the indicator function. By iterating over all voxels and their respective faces, $f_{t}$ is constructed, and the cumulative visibility $F_{t}$ is updated accordingly. Although this method cannot handle all face visibilities, the approximation enables efficient computation of face visibility. Unlike prior works [chen2024gennbv], which consider only $O_{t}$ and $C_{t}$ and thereby treat voxels as points, we treat voxels as cubes to mitigate the information loss caused by approximating voxels as points (see Fig. 1). For details regarding $O_{t}$ and $C_{t}$, please refer to the work [chen2024gennbv].
@@ 3.2 Next-Best-View Hierarchical Network
Training loss functions. The look-at point and viewpoint prediction networks can be jointly trained using the RL reward due to their connection. However, our previous experiments showed no clear benefit from joint training using the RL reward. To simplify the design, we detach the gradient flow between the networks and train the look-at point prediction network with supervised learning using ground-truth targets. The ground truth look-at point $L_{t}^{\text{gt}}$ is computed as the weighted average position of the ground truth uncaptured surface:
@@
The loss of the viewpoint prediction network is the same as the regular RL loss $\mathcal{L}_{\text{RL}}$ which depends on the RL method used, combined with an auxiliary loss:
@@ 4.2 Overall Performance
This section addresses three questions: Q1: Is Hestia’s improvement marginal? Q2: Does the method outperform prior works [jin2023neu, guedon2023macarons, peralta2020next, chen2024gennbv] with and without large-scale training [deitke2024objaverse, Deitke_2023_CVPR]? Q3: Does large-scale training further improve performance? To answer these questions, we benchmark on three datasets [deitke2024objaverse, Deitke_2023_CVPR, wu2023omniobject3d, peralta2020next] comprising 400 diverse shapes, ensuring a comprehensive and fair comparison across methods for the point cloud reconstruction task. Given the large-scale test set, we select three generalizable baselines [jin2023neu, chen2024gennbv, peralta2020next] that do not require test-time optimization, along with one online learning approach [guedon2023macarons] for benchmarking.
@@ 4.4 Translation Robustness
This section demonstrates that Hestia maintains robustness when objects are placed at different positions within the scenes. Tabs. S1, S2 and S3 provide detailed results across different object placement settings. Hestia exhibits less than 1% performance fluctuation, outperforming prior methods on all three datasets. The qualitative results (see Figs. 4 and S6) also show that Hestia’s reconstruction is more robust. These results highlight the effectiveness of the hierarchical structure, which first predicts the look-at point and then determines the capture destination.
@@ 6 Acknowledgements
We thank Yu-Lun Liu for insightful discussions on project direction and Yi-Shan Hung for assistance with paper proofreading and demo video narration.

### v3

@@ 1 Introduction
Multiview-based 3D scene reconstruction [schoenberger2016sfm, yao2018mvsnet, xie2019pix2vox, murez2020atlas, wang2021multi, sayed2022simplerecon, xu2023unifying, wang2024dust3r, duisterhof2024mast3r, pan2024glomap] and novel view synthesis [mildenhall2020nerf, yu2021pixelnerf, yu2022plenoxels, mueller2022instant, zheng2024gpsgaussian, gslrm2024, charatan2024pixelsplat, sabour2024spotlesssplats, flynn2024quark, chen2025mvsplat] have been central topics in computer vision. These methods leverage multiview information to reconstruct high-fidelity scenes. However, data acquisition remains a bottleneck. Most data is collected manually, which is time-consuming and labor-intensive, follows preplanned camera trajectories, or relies on non-active capture systems [xu2023vr, broxton2020immersive, lin2021deep, yoon2020novel, li2022neural, lu2024diva, chen2024x360].
@@ 2 Related Work
Scene-specific next-best-view planners. Next-best-view planners have demonstrated promising results in active 3D reconstruction by predicting the optimal viewpoint for data capture based on the current state. Traditional approaches [monica2018contour, zha1997next, liu2018object, hardouin2020surface, hardouin2020next, dai2020fast, zhou2021fuel] rely on hand-crafted rules to determine the next-best viewpoint. For instance, the method [zha1997next] selected the next-best viewpoint by maximizing a rating function favoring smooth regions, which may overlook fine-grained object details. Instead, the methods [monica2018contour, hardouin2020surface] collected data along boundaries between seen and unseen surfaces to capture finer details, but still require handcrafted parameter tuning for each scene. Another approach [liu2018object] scanned segmented objects sequentially using a predefined object database, reducing the need for handcrafted tuning, but its performance degrades in cluttered environments due to inaccurate object matching. Recent advances in deep learning and increased computational power have given rise to learning-based methods [lin2022active, pan2022activenerf, lee2022uncertainty, zhan2022activermap, jin2023neu, sunderhauf2023density, chen2024gennbv, peralta2020next, ran2023neurar, zeng2022efficient, tao2022seer, wilson2025pop, li2025activesplat]. Some studies [lin2022active, sunderhauf2023density] used NeRF [mildenhall2020nerf] ensembles to estimate uncertainty via model disagreement for viewpoint selection, resulting in linearly increasing computational overhead. Other works [pan2022activenerf, ran2023neurar, zeng2022efficient] avoid the computational overhead by incorporating Bayesian-based NeRF [shen2021stochastic, martin2021nerf] to estimate uncertainty for viewpoint selection. Meanwhile, other approaches [lee2022uncertainty, zhan2022activermap] defined the next-best viewpoint as the viewpoint that maximizes the entropy of the density field along the camera rays. Although these methods have shown outstanding performance in collecting data, they typically require sampling candidate viewpoints. In addition, their reliance on online learning makes them less suitable for real-time applications.
@@
Generalizable next-best-view planners. Unlike the aforementioned online-learning approaches, the generalizable methods [jin2023neu, chen2024gennbv, peralta2020next] avoid the training process for new scenes, thereby enabling faster next-best-view selection. Prediction time is important for real-world tasks where a robot may run out of battery within a few minutes. Among generalizable methods, prior work [jin2023neu] proposed a Bayesian-based NeRF that selects the next-best viewpoint to maximize view variance without additional training. However, it still requires candidate viewpoint sampling, resulting in performance unsuitable for real-time applications. Instead, another line of generalizable next-best-view approaches [chen2024gennbv, peralta2020next] utilized reinforcement learning to learn a next-best-view planner to bypass the need for sampling candidate viewpoints. Prior work [peralta2020next] proposed learning a 3-DoF next-best-view policy using a series of grayscale images as observations. Subsequently, prior work [chen2024gennbv] improved upon this method by incorporating occupancy grids into the observations, which provide explicit geometric information. This enhancement enabled the development of a 5-DoF next-best-view planner, achieving an outstanding coverage ratio for unknown scenes.
@@
As shown in Tab. S4, compared to the scene-specific methods, Hestia does not require candidate viewpoint sampling or inference-time optimization, thereby enabling more flexible viewpoint prediction. Compared to the generalizable method [chen2024gennbv], Hestia treats voxels as cubes rather than points, enabling more comprehensive capture. In addition, Hestia adopts a close-greedy scheme to mitigate spurious correlations, introduces a hierarchical structure to model the action space, and uses a more diverse dataset to maintain robust performance across varying object shapes and positions. Notably, Hestia’s hierarchical structure addresses the challenge of high-dimensional continuous action search in RL-based generalizable next-best-view planning, which differs from traditional methods [monica2018contour, hardouin2020surface, dai2020fast, zhou2021fuel].
@@ 3.1 RL Problem Definition
where $\mathbbm{1}(\cdot)$ is the indicator function. By iterating over all voxels and their respective faces, $f_{t}$ is constructed, and the cumulative visibility $F_{t}$ is updated accordingly. Although this method cannot handle all face visibilities, the approximation enables efficient computation of face visibility. Moreover, non-visible voxel faces simply contribute no reward and therefore do not affect the next-best-view selection. Unlike prior works [chen2024gennbv], which consider only $O_{t}$ and $C_{t}$ and thereby treat voxels as points, we treat voxels as cubes to mitigate the information loss caused by approximating voxels as points (see Fig. 1). For details regarding $O_{t}$ and $C_{t}$, please refer to the work [chen2024gennbv].
@@ 3.2 Next-Best-View Hierarchical Network
Training loss functions. The look-at point and viewpoint prediction networks can be jointly trained using the RL reward due to their connection. However, our previous experiments showed no clear benefit from joint training using the RL reward. To simplify the design, we detach the gradient flow between the networks and train the look-at point prediction network with supervised learning using ground-truth targets. The entire architecture of Hestia in Fig. 2 is trained together without any pretraining on other datasets or tasks. The ground truth look-at point $L_{t}^{\text{gt}}$ is computed as the weighted average position of the ground truth uncaptured surface:
@@
The loss of the viewpoint prediction network is the same as the regular RL loss $\mathcal{L}_{\text{RL}}$ which depends on the RL method used (see Sec. S10), combined with an auxiliary loss:
@@ 4.2 Overall Performance
This section addresses three questions: Q1: Is Hestia’s improvement marginal? Q2: Does the method outperform prior works [jin2023neu, guedon2023macarons, peralta2020next, chen2024gennbv] with and without large-scale training [deitke2024objaverse, Deitke_2023_CVPR]? Q3: Does large-scale training further improve performance? To answer these questions, we benchmark on three datasets [deitke2024objaverse, Deitke_2023_CVPR, wu2023omniobject3d, peralta2020next] comprising 400 diverse shapes, ensuring a comprehensive and fair comparison across methods for the point cloud reconstruction task. Given the large-scale test set, we select three generalizable baselines [jin2023neu, chen2024gennbv, peralta2020next] that do not require test-time optimization, along with one online-learning approach [guedon2023macarons] for benchmarking. We do not include 3DGS-based online-learning methods [wilson2025pop, li2025activesplat, khass2025active] as baselines due to differences in data modality, nor methods [li2025nextbestpath, chen2025gleam] that target non-object-centric scenes with fewer degrees of freedom.
@@ 4.4 Translation Robustness
This section demonstrates that Hestia maintains robustness when objects are placed at different positions within the scenes. Tabs. S1, S2 and S3 provide detailed results across different object placement settings. Hestia exhibits less performance fluctuation, outperforming prior methods on all three datasets. The qualitative results (see Figs. 4 and S6) also show that Hestia’s reconstruction is more robust. These results highlight the effectiveness of the hierarchical structure, which first predicts the look-at point and then determines the capture destination.
@@ 6 Acknowledgements
This work was supported in part by the Australian Research Council (ARC) under discovery grant DP220100803 and DP250103612 and ARC Research Hub for Human-Robot Teaming for Sustainable and Resilient Construction (ITRH) grant IH240100016, and Australian National Health and Medical Research Council (NHMRC) Ideas Grant APP2021183. Research was also sponsored in part by the Australia Advanced Strategic Capabilities Accelerator (ASCA) under Contract No. P18-650825 and ASCA EDT DA ID12994, and the Australian Defence Science Technology Group (DSTG) under Agreement No: 12549. We thank Yu-Lun Liu for submission guidance, Yi-Shan Hung for proofreading and demo narration, Meredith Porte for additional narration, and Xiao Chen for GenNBV [chen2024gennbv] discussions.